# Tokens for Learning, Tokens for Unlearning: Mitigating Membership Inference Attacks in Large Language Models via Dual-Purpose Training

### v1

@@ 5.2 Overall Evaluation
Figure 3: Privacy-utility trade-off of the methods while varying hyper-parameters. The Goldfish masking rate is set to 25%, 33%, and 50%. The privacy budget $\epsilon$ of DPSGD is evaluated at 8, 16, 50, and 100. The weight $\alpha$ of DuoLearn is configured at 0.2, 0.5, and 0.8.
@@ 5.4 Training Dynamics
Memorization and Generalization Dynamics. Figure 6 (left) illustrates the training dynamics of conventional fine tuning and DuoLearn, while Figure 6 (middle) depicts the membership inference risks. Generally, the gap between training and testing loss of conventional fine-tuning steadily increases overtime, leading to model overfitting and high privacy risks. In contrast, DuoLearn maintains a stable equilibrium where the gap remains more than 10 times smaller. This equilibrium arises from the dual-purpose loss, which balances learning on hard tokens while actively unlearning memorized tokens. By preventing excessive memorization, DuoLearn mitigates membership inference risks and enhances generalization.
@@
Gradient Conflicts. To study the conflict between the learning and unlearning objectives in our dual-purpose loss function, we compute the gradient for each objective separately. We then calculate the cosine similarity of these two gradients. Figure 6 (right) provides the cosine similarity between two gradients over time. During training, the cosine similarity typically ranges from -0.15 to 0.15. This indicates a mix of mild conflicts and near-orthogonal updates. On average, it decreases from 0.05 to -0.1. This trend reflects increasing gradient misalignment. Early in training, the model may not have strongly learned or memorized specific tokens, so the conflicts are weaker. Overtime, as the model learns more and memorization grows, the divergence between hard and memorized tokens increases, making the gradients less aligned. This gradient conflict is the root of the small degradation of language modeling performance of DuoLearn compared to the conventional fine tuning approach.
Token Selection Dynamics. Figure 7 illustrates the token selection dynamics of DuoLearn during training. The figure shows that the token selection process is dynamic and changes over epochs. In particular, some tokens are selected as an unlearning from the beginning to the end of the training. This indicates that a token, even without being selected as a learning token initially, can be learned and memorized through the connections with other tokens. This also confirms that simple masking as in Goldfish is not sufficient to protect against MIAs. Additionally, there are a significant number of tokens that are selected for learning in the early epochs but unlearned in the later epochs. This indicates that the model learned tokens and then memorized them over epochs, and the during-training unlearning process is essential to mitigate the memorization risks.
@@ 6 Conclusion
We introduced DuoLearn, an effective training framework defending against MIAs for LLMs. The extensive experiments demonstrate its robustness in protecting privacy while maintaining strong language modeling performance across various datasets and architectures. Although our study focuses on fine-tuning due to computational constraints, DuoLearn can be seamlessly applied to large-scale pretraining, as done in prior selective pretraining work Lin et al. (2024). By categorizing tokens and treating them appropriately, DuoLearn opens a novel pathway for MIA defense. Future work can explore improved token selection strategies and multi-objective training approaches.
@@ Appendix B Token-level membership inference risk analysis
Figures 8 and 9 present the analysis for additional samples. Generally, the trends are consistent with the one presented in Section 3.
@@ D.2 Auxiliary dataset
D.3 Training time
@@
Appendix E Limitations
The main limitation of our work is the small-scale experiment setting due to the limited computing resources. However, we believe DuoLearn can be directly applied to large-scale pretraining without requiring any modifications, as done in previous selective pretraining work Lin et al. (2024). Another limitation is the reference model, which may be restrictive in highly sensitive or domain-limited settings Tramèr et al. (2024). From a technical perspective, while we show that DuoLearn performs well across different datasets and architectures, there is room for improvement. The current approach selects a fixed number of tokens, which may not be optimal since selected tokens contribute unequally. Future work could explore adaptive selection or weighted tokens’ contribution. At a high-level, compared to DPSGD, DuoLearn has not been supported by theoretical guarantees. Future work can investigate the convergence and overfitting analysis.

### v2

@@ 5.2 Overall Evaluation
Figure 3: Privacy-utility trade-off of the methods while varying hyper-parameters. The Goldfish masking rate is set to 25%, 33%, and 50%. The privacy budget $\epsilon$ of DPSGD is evaluated at 8, 16, 50, and 100. The weight $\alpha$ of DuoLearn is varied at 0.0, 0.1, 0.2, 0.5, and 0.8.
@@ 5.4 Training Dynamics
Memorization and Generalization Dynamics. Figure 6 (left) illustrates the training dynamics of conventional fine-tuning and DuoLearn, while Figure 6 (middle) depicts the membership inference risks. Generally, the gap between training and testing loss of conventional fine-tuning steadily increases over time, leading to model overfitting and high privacy risks. In contrast, DuoLearn maintains a stable equilibrium where the gap remains more than 10 times smaller. This equilibrium arises from the dual-purpose loss, which balances learning on hard tokens while actively unlearning memorized tokens. By preventing excessive memorization, DuoLearn mitigates membership inference risks and enhances generalization.
@@
Gradient Conflicts. To study the conflict between the learning and unlearning objectives in our dual-purpose loss function, we compute the gradient for each objective separately. We then calculate the cosine similarity of these two gradients. Figure 6 (right) provides the cosine similarity between two gradients over time. During training, the cosine similarity typically ranges from -0.15 to 0.15. This indicates a mix of mild conflicts and near-orthogonal updates. On average, it decreases from 0.05 to -0.1. This trend reflects increasing gradient misalignment. Early in training, the model may not have strongly learned or memorized specific tokens, so the conflicts are weaker. Overtime, as the model learns more and memorization grows, the divergence between hard and memorized tokens increases, making the gradients less aligned. This gradient conflict is the root of the small degradation of language modeling performance of DuoLearn compared to the conventional fine-tuning approach.
Token Selection Dynamics. Figure 7 illustrates the token selection dynamics of DuoLearn during training. The figure shows that the token selection process is dynamic and changes over epochs. In particular, some tokens are selected for unlearning from the beginning to the end of the training. This indicates that a token, even without being selected as a learning token initially, can be learned and memorized through the connections with other tokens. This also explains that simple masking as in Goldfish is not sufficient to protect against MIAs. Additionally, there are a significant number of tokens that are selected for learning in the early epochs but selected for unlearning in the later epochs. This indicates that the model gradually memorizes these tokens over epochs, and the during-training unlearning process is essential to mitigate the memorization risks.
@@ 5.5 Privacy Backdoor
Method Eval Loss (↓) MIA AUC (↓) TPR @ 1% FPR (↓) Training Time (↓)
Untrained model 11.329 0.517 0.010 N/A
Conventional LM 3.003 0.930 0.376 10.095 hours
Goldfish 2.844 0.905 0.426 10.159 hours
DuoLearn 3.244 0.548 0.040 12.278 hours
Table 3: Comparison of Methods on Evaluation Loss, Privacy Metrics, and Training Time for pretraining.
5.6 Pretraining
We conduct a small-scale pretraining experiment using a Llama-like architecture with 1.5 billion parameters. The experiment is to pretrain on a dataset of 1 billion tokens. We reuse the dataset and codebase developed by Sanyal et al. (2024) [^4] Following this setting and due to limited computing resources, we use a batch size of 8 which is much smaller than practical pretraining. All methods implement the same learning rate and are evaluated at their 25K-th iteration.. The pretraining corpus is collected from various sources and domains, including arXiv, books, Common Crawl, GitHub, StackExchange, and Wikipedia Weber et al. (2024). To train the reference model, we use 10% of the data. Table 3 presents the performance of DuoLearn in this pretraining. Generally, DuoLearn successfully mitigates the MIA risk, reducing AUC from 0.9 to 0.55 and TPR at 1% FPR from 0.4 to 0.04, with minimal degradation on the model performance.
5.7 Low-Prevalence Tokens vs High-Prevalence Tokens
To understand the bias of the methods towards low-prevalence and high-prevalence tokens, we conduct some visualization on the token-level MIA signals and token frequencies. We selected a set of 500 samples in the training data. We split the tokens into two sets, high-frequency and low-frequency, then visualize kernel density estimation (KDE) separately, illustrated in Figure 8. Notably, both conventional FT and DuoLearn yield different privacy risks for the two groups. The low-prevalence tokens have higher per-token MIA risks (with MIA signal ranging from -5 to 2, compared to -0.5 to 0.5 of high-prevalence tokens). DuoLearn successfully shifts the MIA-signal distribution to be centered at zero for both high-frequency and low-frequency tokens. This indicates that DuoLearn reduces the majority of MIA risks across both token groups. However, it slightly increases the variance of risks for both groups.
Figure 8: Effects on high-prevalence and low-prevalence tokens. For MIA signal, we consider the Ref-Loss attack, a signal value closer to zero indicates lower risk.
@@ 6 Conclusion
We introduced DuoLearn, an effective training framework defending against MIAs for LLMs. The extensive experiments demonstrate its robustness in protecting privacy while maintaining strong language modeling performance across various datasets and architectures. Although our study focuses on fine-tuning and small scale pretraining due to computational constraints, DuoLearn can be seamlessly applied to large-scale pretraining, as in prior selective pretraining work Lin et al. (2024). By categorizing tokens and treating them appropriately, DuoLearn opens a novel pathway for MIA defense. Future work can explore improved token selection strategies and multi-objective training approaches.
Limitations
The main limitation of our work is the small-scale experiment settings due to limited computing resources. However, we believe DuoLearn can be directly applied to large-scale pretraining without any modifications, as in previous reference-model-based pretraining study Lin et al. (2024). Another limitation is the reference model, which may be restrictive in highly sensitive or domain-limited settings Tramèr et al. (2024). From a technical perspective, while DuoLearn performs well across different datasets and architectures, there is room for improvement. For example, future work could explore adaptive selection size or weighted token contribution. Additionally, as DuoLearn is an empirical defense, future work can investigate the convergence and overfitting analysis.
Acknowledgment
This work is partially supported by the National Science Foundation under Award Numbers 2302968, 2124104, and 2125530, and by the National Institutes of Health under Award Numbers R01ES033241 and R01LM013712. The views and opinions expressed in this paper are those of the authors and do not necessarily reflect the views of the U.S. Government or any agency thereof.
@@ Appendix B Token-level membership inference risk analysis
Figures 9 and 10 present the analysis for additional samples. Generally, the trends are consistent with the one presented in Section 3.
@@ D.2 Auxiliary dataset
D.3 Hyperparameter sensitivity analysis
$\bullet$ Varying $\mathcal{T}_{h}$ and $\mathcal{T}_{m}$ – Portion of tokens for learning and for unlearning. We keep other hyperparamters as default and adjust $\mathcal{T}_{h}$ and $\mathcal{T}_{m}$ separately. Figures 13 and 14 provides the results of this experiment. Generally, DuoLearn is robust while varying these hyperparameters with PPL ranging from 22 to 23 and AUC ranging from 0.46 to 0.56.
Figure 13: Performance of DuoLearn while varying $\mathcal{T}_{h}$
Figure 14: Performance of DuoLearn while varying $\mathcal{T}_{m}$
$\bullet$ Varying $\alpha$ – Weight balance factor. Figure 15 illustrates the peformance while varying $\alpha$. Intuitively, the smaller $\alpha$, the less unlearining performed, it leads to a better language modeling performance and higher privacy risk. When $\alpha$ is unreasonably high (i.e., 1.5 or 2.0), the unlearning part dominates the learning one, it leads to high perplexity values of language modeling.
Figure 15: Performance of DuoLearn while varying $\alpha$
D.4 Training time